LUCIFER Com2
UT 2008 December 08

Observers: N.Ageorges, W.Seifert, M.Juette, F.Eisenhauer, D.Thompson
Telescope Operator: D.Gonzalez-Huerta
SW Support: T.Edgin, on call

Data will eventually be copied to /Repository/LUCIFER_Data/20081205/

Summary:

Details:

D.Gonzalez-Huerta:
  • 15:20 Started walk-through, telescope checkout.
  • 15:30 LBC Blue filled with LN2
  • 16:00 Test preset completed successfully.
  • Encoders reset (worked the first time)
  • 16:07 AGW motors homed (not stopped or restarted)
  • 16:12:29 Telescope at zenith, rotator at 341, turned over to LUCIFER Team

16:14 Weather: outside temp 0.7C, RH 81% and rising, 10m/s wind (SMT 1m/s). Currently completely cloudy, occasional snow this afternoon.

We remained closed the entire night. LUCIFER Team continued with calibration unit, working on flexure compensation and other things.

Astrometric data from a previous night shows no trends with time or EL, and rms variations in scale throughout the night are at the 0.028% level. There is marginal evidence for slightly larger variations from one night to the next (0.04%), but more analysis of exisiting data is needed.

-- DavidThompson - 05 Dec 2008
Topic revision: r3 - 11 Dec 2008, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback